OpenStudy (anonymous):

What is 3/10 divided by 4/5 in simplest form?

OpenStudy (anonymous):

3/8

OpenStudy (anonymous):

3/8

OpenStudy (anonymous):

3/8

OpenStudy (anonymous):

or 0.375

OpenStudy (paki):

(3/10) / (4/5) = 15/40 = 3/8

OpenStudy (anonymous):

Because when you divid, your really mult. so you would flip the 4/5 to 5/4 then do the math, and you get15/40 then use simpest form and get 3/8.